Engineering/D-Night 20180505 (UT)

Night Plan:

Before Opening/Twilight:
  • DONE AO Startup with OSA (Doug, Al, David G.H.)
  • DONE Verification of Conditions (Doug, Al, David G.H.)
    • Doug has scripts that collect telemetry and the 3 representative LUCI images. The plan is to try these on sky.
    • We'll use IDL for the runAO etc since IRAF mods not yet ready (to the best of our knowledge)
  • Total Time for AO startup activities ~ 1.5 hours
After Opening/Twilight:
  • DONE Quasi-Earth Satellite Observations - Binocular LUCI - zJ G200 spectroscopy (Al, Olga, Barry) ~ 2 hours
    • Seeing Conditions - up to 1.2"
    • Non-sidereal
    • V ~ 22
  • DONE AO LookUp Table testing (Doug M. responsible for Scripts/Data collection) ~ 1.5 hours - TOP PRIORITY
    • Requires Seeing </ 1.2"
    • Can be done with some extinction (may be useful to cut down flux for bright Bin 1 targets - should be constant)
  • DONE AO/ESM/SL comparison of the same target ~ 1.5 hours
    • DONE Potential Target 1: M92 - offset from center - 17:17:07 +43:09:09
    • Potential Target 2: SH106 - star cluster - end of night 20:27:25 +37:22:45
    • Should be done with seeing < 1.2"
    • UPDATES TO TCS/GCS:
      • updated GCS Version 5.6a (Kellee) for the upcoming E-night. Two simple fixes - these do not need specific on-sky tests, can be tested while other observations underway
        • IT7206 - offset distance in log message should use binned pixelscale
        • IT7045 - centroiding sometimes gets hung and crashes GCS
      • TCS work (Petr) -these do not need specific on-sky tests, can be tested while other observations underway
        • fixed IT7218 (NS DETXY breaks co-pointing limits; this does not fix RADEC cause, so my work of cautions to Al&co about disabling co-pointing checks while running NS observations is still valid)
        • refactoring of TrajectoryGenerator code in PCS to support calculations of DX tip/tilts for co-pointing checks (J.Hill is aware of this)
        • PCS will not generate events with wrong ephemerides (too old ephemeris) at 2 Hz, will run at 1/60 Hz (=every 1 minute) instead
        • minor fixes in ephemeris parsing code (so wrongly formatted, e.g. missing $EOE$ file, will be properly handled)
        • upgrades of ephemeris file reading locks
After Opening/Twilight if conditions include seeing > 1.2"-1.5" and/or significant extinction from clouds:
  • Quasi-Earth Satellite Observations - Binocular LUCI - zJ G200 spectroscopy (Al, Olga, Barry) ~ 2 hours
    • Seeing Conditions - up to ?
    • Non-sidereal
    • Some extinction?
  • Non-Linearity on-sky tests (Dave T.) ~ 30-45 min
  • Barry - Near-IR zJ,HK observations of Local Mergers
    • Seeing up to ~2" - extinction up to 3-4 mags
    • NGC 4441

      Telluric HD 116405

      12 27 20

      13 22 45.12

      64 48 06

      44 42 53.9

      Arp 156

      Telluric AG+69 313

      10 42 38

      06:47:28.808

      77 29 41

      69:37:44.77

      NGC 2655

      Telluric AG+69 313

      08 55 37

      06 47 28.81

      78 13 23

      69 37 44.8

      NGC 3310

      Telluric HD 92573

      10 38 45

      10 42 7.67

      53 30 05

      45 50 54.2

      UGC 4079

      Telluric HD 56385

      07 55 06

      07 19 51.28

      55 42 13

      54 55 18.1

      NGC 2744

      Telluric BD+21 1974

      09 04 38

      09 07 55.69

      18 27 37

      20 53 49.6

      NGC 6598

      Telluric HIP 82884

      18 08 56

      16:56:18

      69 04 04

      52:41:54

      NGC 4004

      Telluric HD 109055

      11 58 05

      12 31 41.28

      27 52 44

      22 7 24.4

  • Barry - H-band Library (seeing >>2" >>3-4 mags extinction)
  • Olga -
    • NGC 4150 12:10:34 +30:24:06 (Telluric 12:36:24 +32:11:45)
    • Q1330+011


Night Log

AO Startup/Checkout (E-time)

OSA: D.G. Huerta ISA: B. Rothberg Observers: B. Rothberg, O. Kuhn ISp: C. Vogel

UT 00:06 - Taking LUCI-Bino Darks for AZ Observers

UT 01:00 - Darks complete

UT 01:49 - starting AO checkout with Doug, Al, and David G.H.

UT 02:08 - Enclosure Opening

UT 02:31 - configuring LUCIs for test exposures, N30, N30fieldstop, OH1060+z (will not give representative image of AO corrections though)

UT 02:54 - starting up AO...

UT 02:58 - some issues with SX AO - raided leftover food from all-hands meeting to compensate... (seeing at start 2.5" DIMM now reports 1.5" guiders report 1.2")

UT 03:02 - mirror panic on SX - preset cancelled

UT 03:09 - SX primary back up - SX AO is having communications errors - Doug M. suggested we finish the AO startup, will recover this while switch to D-time for the Quasi-satellite

Quasi-Earth Satellite (Mostly D-Time/)

UT 03:14 - doing pointing and collimation check for SL for science target

UT 03:32 - fixed collimation error (runaway z11/z22 on DX)

UT 03:33 - sending preset to Quasi - NSIGUI hijack....hit co-pointing limit - this may be caused by the 1.5/2 slit because it has an iniital 47" shift (along with NS) to get you to the 1.5" slit - although the 47" movement is supposed to by a mount movement (sync offset)

UT 03:37 - resending preset

UT 03:39 - LUCI-1 MOS Error - error returning mask 4 (1.5"/2.0" longslit) to storage - mask had been in FPU - called Dave T. - turned off mirror,instrument, and readout for LUCI-1

UT 03:49 - -skipped to 2nd line in script, then clicked "continue with LUCI-2" - field image being taken - got a longslit image - but this is because Barry put mask in FPU earlier. Moving mask to FPU to take our 2nd acq image. L2#0039 bkgs, L2#0040, L2#0041 - field image - seems we can proceed with L2 while L1 is out of commission.

UT 03:58 - guider DX 0.85" avg fwhm

UT 04:05 - Dave T. recovered error on LUCI-1 MOS, init IMGUI - will resend preset - increasing exposure time on acquisition images, didn't see target in 120sec

UT 04:09 - LUCI-1 MOS error is related mask slot 4 - the open switch is red - this also caused an issue during the last exchange - also turned off check co-pointing switch in PCS GUI

UT 04:15 - PCS error - offset - Petr K. fixing it -

UT 04:16 - recovering from error, will need to send preset again - preset is not active although system is not tracking

UT 04:18 - sending preset (NSIGUI hijacked)

UT 04:20 - hotspot on guiders seem off to the right... but its working guider shows ~ 0.5" on both sides

UT 04:39 - doing longslit acq - Barry on LUCI-2 Olga on LUCI-1 - guider seeing 0.56" 0.57"

UT 04:50 - starting spectroscopic observation - left in RA Dec absorb in script - but didn't lose preset this time

UT 04:55 - PCS died after first exposure, preset died

UT 04:57 - reverting PCS to old version - Olga will edit script to remove the RA Dec absorb

UT 05:10 - resending preset - how to edit the script for reduced exptimes in acq images and thru-slit, had an error where DX side failed to find guidestar and complained of moving AGW too far, however, actual positions were very close on both sides.

UT 05:13 - sent script on LUCI, but telescope doesn't seem to be getting the preset - problem was the weird state it gets into sometimes that even though preset fails, it hangs on waiting for collimation.

UT 05:14 - third attempt preset went through

UT 05:25 - can see the object on both sides in acq images in 3minutes

UT 05:33 - wfs still shows paused in guider, but the hotspot is not jumping around like it did for the PCS update

UT 05:37 - spectroscopic integration started

UT 05:53 - WFS on SX last corrected 05:33 - looks like a trianlge in the guidecamera

UT 06:09 - PCS died - reverting PCS did not seem to fix it, but we did last longer (1 exposure before)

UT 06:10 - going to nearby star to point - put in mirror, blind filters, and turned off flexure

UT 06:19 - pointied and collinated (since WFS doesn't seem to update) - doing updated script

UT 06:26 - changing wfs exposure to 5sec

UT 06:40 - DIMM showing 1.3, but guiders <1"

UT 06:41 - starting spectroscopic exposures

UT 06:56 - DX is now making active optics corrections (starting 6 min ago), but SX is not.

UT 07:03 - manual focus added on SX

UT 07:09 - hard to tell if see continuum, SX still not wavefront sensing

UT 07:15 - moving to take Telluric star spectroscopy

UT 07:23 - not hitting the guider hotspot - not enough flux on SH spots

UT 07:28 - moving to the solar analog - the telluric (which is actually a G0) appeared to be a double (SIMBAD notes its high proper motion and shows no double so the other star could be a background one due to high proper motion)

UT 07:30 - pointing error - guidestar failed one one side - doing pointing check

UT 07:34 - sending preset to solar analog

UT 07:36 - collimation is terrible - perhaps it is seeing.....

UT 07:43 - guiders showing 1.5" and 1.6"

UT 07:55 - DIMM 1.65"

Last L1#0074, Last L2#0077

AO-Testing/Comparison Imaging (E-time)

UT 07:56 - getting AO ready - moving to zenith - LUCIs in safe mode while doing this....

UT 07:59 - AO ready, doing pointing and collimation checks

UT 08:10 - Sending ESM preset to M92

UT 08:14 - 1.49" on the DIMM

UT 08:19 - SX wonky, will continue taking data - this will be the SL side (while Doug works on it), DX will be the ESM (so getting simultaneous data)

UT 08:24 - still waiting for AO to go ahead....

UT 08:26 - Left side failed and would not continue, Doug is sending an active (not adaptive) on the left to see if we can continue

UT 08:27 - AO system is waiting on DX - in a funky state.

UT 08:27 - Doug M. will send a resume - DX still waiting

UT 08:28 - Doug M. sending a preset on left to see if this will bump it to continue - didn't work

UT 08:30 - John H. and Doug M. killed offset, need an ACTIVE offset on left

UT 08:34 - Left failed because its expecting an AO

UT 08:36 - Doug M. will restart left, then we will resend the preset and continue in bino (DX attempted offset and Doug M. lost the star)

UT 08:37 - sending binocular preset

UT 08:40 - DX preset failed - AO exception (apparently this type of error happens once in a while) - Doug M. working on it.

UT 08:41 - resent preset - DIMM 2.5"

UT 08:42 - did the new preset arrive? Doug M. and John H. don't think so

UT 08:43 - sending preset again

UT 08:50 - seems to be wokring, but AO reports R=12.8, star is R=9.4

UT 08:52 - Dimm reports 1.3", FWHM of star is 1" with ESM

UT 09:00 - after return from off fields, DX caught the star, SX the star flew by - Doug M. recovering

UT 09:03 - skipped to same line in script, sent GO....loops closed on both sides...

UT 09:04 - DIMM 1.18" - ESM getting ~0.5" on both sides (Bin 4 400 Hz)

UT 09:14 - loops opened - are paused

UT 09:18 - loops closed - continuing...

UT 09:29 - AO finished, sending SL script that does the same thing - guidestars different on each side (were fine for ESM, but somehow pointing degraded and ~15" offset in guidestars)

- L1#0075-0089 TT - L2#0078-0092 ESM images

UT 09:32 - accidentally sent a guide preset not active - fixed script

UT 09:37 - resent picked a guidestar (OT seemed to have deleted the GS and selected the AOref as GS) - but forgot to fix guide mode

UT 09:39 - resending active preset (L1#092 start of SL, L2#0095 start of SL) - guider showing ~1.2", DIMM 1.7"

UT 10:04 - DIMM 1.8-1.9" ~1.2" on guiders L1#0092-0106, L2#0095-0109 SL

UT 10:07 - starting fainter ESM test (R=14.6 on UCAC5 - update rate 200Hz Bin 4 - R=14.7 reported on wavefront sensor - looks like test of extreme range) DIMM 1.5"

UT 10:30 - attempted to recover by loading script again and going to line 2, IIF choked and entire thing failed

UT 10:32 - turning off right side, trying to continue left only - didn't work - taking a manual exposure on the left - didn't work - seems to hang had to abort

- aborted something and then left side kicked in and started exposure, except 'scope moved and star is not in AO system -didn't save the file -

TTLUT (E-Time)

UT 10:41 - configuring LUCIs for Doug M. HKspec+clear, N30fieldstop - Camera on LUCI-1 unknown

UT 10:44 - running init on LUCI-1 Camera - took a bit of time, realized it was on N375, configuring - 'scope moving to SE

UT 10:49 - DIMM reports 2.8"

UT 10:58 - moved to new target - guiders show ~1.5"

UT 11:06 - aligning N30 fieldstop on both sides

UT 11:13 - Doug starting his script, should take exposures every 60seconds

UT 12:02 - loop opened on SX and DX - taking darks on each

UT 12:06 - done, but Doug M. is having an issue on the right side trying to resolve

UT 12:12 - closing up, will need to take cals (K imaging flats N375 and 1" LS)

Closed Dome Calibrations

UT 12:34 - Starting K N3.75 calibrations (flats)

L1 #0145-0154

L2 #0148-0157

UT 12:29 - Starting 1"LS G200 zJ calibrations

-- %USERSIG{BarryRothberg - 2018-05-03}%

Topic revision: r6 - 07 May 2018, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback